the long-distance calls made by employees of a company are normally distributed with a mean of 6.3 minutes and a standard deviation of 2.2 minutes. Find the probability that a call
a. last between 5 and 10 minutes
b. last more than 7 minutes
c. last less than 4 minutes
1
Expert's answer
2019-09-23T09:24:48-0400

Let "X" be a random variable which denotes the call duration in minutes. Our random variable "X" is normal

"X\sim N(\mu,\sigma^2)". Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=6.3 \ minutes, \sigma=2.2 \ minutes."

a.


"P(5<X<10)=P\bigg({5-6.3 \over 2.2}<Z<{10-6.3 \over 2.2}\bigg)=""=P\bigg(Z<{3.7 \over 2.2}\bigg)-P\bigg(Z<{-1.3 \over 2.2}\bigg)\approx""\approx0.953698-0.277291\approx0.6764"

b.


"P(X>7)=P\bigg(Z>{7-6.3 \over 2.2}\bigg)=1-P\bigg(Z<{0.7\over 2.2}\bigg)\approx""1-0.624826\approx0.3752"

c.


"P(X<4)=P\bigg(Z<{4-6.3\over 2.2}\bigg)=P\bigg(Z<{-2.3\over 2.2}\bigg)\approx""\approx0.1479"